OpenStudy (harlequin):

(5x-2)^1/3=2 How do I work his out I have the answer but I don't know how to get it.

OpenStudy (_malice_):

What do you have for the answer?

OpenStudy (harlequin):

The answer end up being two

alones (alones):

\[(5x-2)^\frac{ 1 }{ 3 }=2\] Like this?

OpenStudy (harlequin):

Yep

OpenStudy (_malice_):

Solve for x by simplifying both sides of the equation, then isolating the variable. right alone bb?

alones (alones):

Oh well yes 2 is right..let me explain this

alones (alones):

Yes correct malice bb (:<

OpenStudy (_malice_):

;D I AM THE SMERT

alones (alones):

Solve Exponent \[(5x−2)^\frac{ 1 }{ 3 }=2\] \(\ ((5x−2)^\frac{ 1 }{ 3 })^3=2^3\)<--Raise both sides to power \(\ 5x−2=8\)

OpenStudy (_malice_):

im gonna go, bubai alone bb <3

alones (alones):

Than add those two sides \(\ 5x−2+2=8+2\) \(\ 5x=10\) and afterall divide both sides by 5 \[\frac{ 5x }{ 5 }=\frac{ 10 }{ 5 }\] \[\ x=2\]

alones (alones):

@Harlequin Now you understand?

OpenStudy (harlequin):

Thank you now I now how to do the rest of the chapter.

alones (alones):

You welcome hun

OpenStudy (harlequin):

Yes, I took notes but the instructor didn't go over this part of the lesson so I had no idea what I was doing. The examples in the book didn't go over it either so this helped a lot.